# Strada le Grazie 15, Verona, Italy

### v1

@@ 4.3 Varying Noise Levels on KSDD2
In this study, we rigorously evaluate the robustness and efficacy of the Iterative Refinement Process (IRP) on the KSDD2 dataset. Our analysis delves into the performance dynamics of the IRP model under various noise conditions, revealing its exceptional defect detection capabilities in the face of escalating environmental challenges. As illustrated in Fig. 3, the IRP model is compared against two conventional approaches—the traditional vanilla model and the One Shot Removal (OSR) model proposed in [27]. Across all tested noise levels, the IRP model consistently outperforms the competitors, demonstrating remarkable resilience by maintaining superior AUROC values, as systematically documented in Table 1.
The IRP model’s prowess is particularly pronounced at elevated noise levels. For instance, at a substantial noise threshold of 50%, the IRP model sustains a significantly higher AUROC of 0.8967 on the KSDD2 dataset. This performance starkly contrasts with the pronounced declines observed in the vanilla and OSR models. This robustness is not merely a result of the algorithm’s inherent strength but arises from the IRP’s comprehensive application of advanced noise-cancellation algorithms, adaptive thresholding techniques, and its capacity to effectively extract complex features from heavily corrupted data streams.
@@
Moreover, the statistical analysis accompanying our findings reinforces the IRP model’s superiority. The standard deviation values reported in Table 1 indicate lower variability in performance across different experimental runs, underscoring the model’s reliability and predictability. Such attributes are indispensable in industrial applications, where the cost of false negatives can be prohibitive, and the ability to detect anomalies reliably is paramount. The comprehensive data presented in the table not only showcases the IRP model’s adaptability and robust performance across diverse and challenging scenarios but also highlights its viability as a potent tool for maintaining stringent quality control standards in manufacturing processes. These findings suggest that the IRP can be effectively deployed in industrial settings to enhance defect detection where traditional models might falter due to high variability in defect rates and challenging noise conditions.
4.4 Varying Noise Levels on MVTec-AD
The robustness and adaptability of IRP have been rigorously evaluated across various product classes within the MVTec-AD dataset, including Cable, Capsule, MetalNut, Pill, Screw, and Zipper. Each class, with its distinct defect types and unique challenges, served as an ideal testbed for the IRP. We systematically varied the noise levels in the dataset from 0% to 50% to meticulously assess the IRP’s capabilities under progressively challenging conditions, using the Area Under the Curve of the Receiver Operating Characteristic (AUROC) to quantify defect detection accuracy across different thresholds.
As illustrated in Fig. 4, the IRP model consistently achieves high AUROC values across all classes, demonstrating its robust ability to manage noise effectively. Notably, in classes with complex and variable defect characteristics such as Screw and Cable, the IRP exhibits exceptional resilience, showcasing its advanced anomaly detection capabilities. Conversely, in classes like Capsule and Pill, where defects are more subtle, the IRP effectively distinguished between normal variations and actual defects, underlining its sophisticated feature extraction capabilities. This uniform performance across varied classes and noise levels highlights the IRP model’s versatility and effectiveness in industrial applications, especially in settings where defect rates are unpredictable and varied. The IRP’s ability to maintain high accuracy at elevated noise levels underscores its potential as a dependable tool for quality control in manufacturing environments.
Furthermore, Fig. 3 provides an overarching view of the IRP’s performance across the MVTec-AD dataset. This visual representation emphasizes the IRP’s superior performance relative to the vanilla and OSR models, particularly as it consistently improves defect handling under increasingly challenging conditions. This enhancement affirms the IRP’s role as a robust and reliable quality control instrument in complex manufacturing landscapes, where defect rates can fluctuate significantly. Table 1 complements these insights by detailing the AUROC scores at various noise levels for both the MVTec-AD and KSDD2 datasets, including measures of error magnitude. This detailed presentation not only confirms the performance stability of the IRP across different settings but also highlights the model’s reliability, evidenced by the low standard deviation in its performance metrics.
@@ 4.4 Varying Noise Levels on MVTec-AD
Noise Level (%) Capsule Cable MetalNut Pill Screw Zipper MVTec-AD KSDD2 Error Magnitude
0% 0.8592 0.9288 0.8880 0.8888 0.8938 0.9378 0.8994 0.9403 ±0.0246
10% 0.8498 0.8368 0.8688 0.8580 0.868 0.8880 0.8615 0.9238 ±0.0149
20% 0.8288 0.856 0.8480 0.8558 0.8524 0.8504 0.8485 0.9024 ±0.0086
30% 0.8102 0.6756 0.8464 0.8016 0.8234 0.8232 0.7967 0.9109 ±0.0518
40% 0.8012 0.5716 0.7792 0.8096 0.7988 0.8096 0.7616 0.8923 ±0.0777
50% 0.7872 0.4604 0.8096 0.7816 0.7436 0.7340 0.7193 0.8967 ±0.1099
@@
4.5 Quantitative Analysis of Removed Samples
In this section, we present a detailed quantitative analysis of the samples removed during the defect detection process for both the KSDD2 and MVTec-AD datasets. This analysis is segmented by product categories and noise levels, providing insights into the effectiveness of the Iterative Refinement Process (IRP) in improving defect detection accuracy. Table 2 shows the number of good and bad samples removed at various noise levels (0% to 50%) across different product categories within the KSDD2 and MVTec-AD datasets. The categories include Cable, Capsule, Metalnut, Pill, Screw, and Zipper. Each entry in the table specifies the count of good (non-defective) and bad (defective) samples removed during the training and validation phases. The accompanying graph Fig. 5 provides a visual representation of the number of good and bad samples removed at each noise level, helping to understand the distribution and trends of sample removal across different noise conditions.
@@ 4.5 Quantitative Analysis of Removed Samples
The table indicates that at 0% noise, only good samples are removed. This might seem counterintuitive, but it can be explained by the fact that these good samples are likely close to the decision boundary, making them appear anomalous during the initial iterations of the IRP. Removing these borderline good samples helps in refining the model by reducing potential noise and improving its generalization capability. This phenomenon demonstrates the IRP’s ability to fine-tune the model even when no actual bad samples are present. As noise levels increase, the number of bad samples removed rises, which is expected. However, the table also shows that good samples continue to be removed across all noise levels. This ongoing removal of good samples close to the boundary helps in maintaining the robustness of the model by continuously refining its decision boundaries.
The quantitative analysis of removed samples underscores the IRP model’s capacity to enhance defect detection accuracy through systematic refinement. By removing samples close to the decision boundary, the IRP improves the overall performance, even in the absence of bad samples. As noise levels increase, the model adeptly identifies and eliminates defective samples while maintaining a balance by also removing borderline good samples. This approach not only fine-tunes the model but also ensures its robustness and reliability in real-world industrial applications.
Figure 5: This graph shows the number of good and bad samples removed at various noise levels, illustrating the distribution and trends in sample removal under different noise conditions.
@@ 5 Discussion
The robust performance of the IRP across varying noise levels underscores its potential utility in industrial settings, where defect patterns may be unpredictable. The model’s ability to maintain high AUROC values under increasing noise conditions illustrates its capacity to effectively distinguish subtle defect features from normal variability an essential attribute for critical quality control environments.
@@
While the model demonstrates strong performance, there is scope to enhance its operational efficiency. The current training setup, designed for moderate scale and tailored to specific experimental conditions, serves well under controlled scenarios. However, as we aim to deploy the model in more dynamic industrial settings, optimizing its capacity to handle larger datasets and faster processing cycles will be crucial. These enhancements are intended to reduce the time required to reach model convergence and increase the throughput of image processing, ensuring the model can meet the demands of real-time defect detection. Future developments could include refining the training algorithms to accelerate convergence without compromising accuracy, and expanding the system’s architecture to support more simultaneous operations. These improvements will be critical for deploying the model in environments where rapid decision-making is essential. Furthermore, the detailed analysis of different classes within the MVTec-AD dataset highlights the IRP’s adaptability to diverse manufacturing scenarios. The pronounced resilience in classes with intricate defect patterns, such as Screws and Cables, underscores the model’s sophisticated feature recognition capabilities, which are likely bolstered by advanced data preprocessing and anomaly scoring mechanisms. Moving forward, additional research could focus on optimizing the model to enhance detection accuracy in scenarios with extremely subtle defect signals, such as those found in high-grade pharmaceutical manufacturing.
@@
Future research efforts might also refine the probabilistic models used within the IRP, enhancing their sensitivity to minor anomalies and incorporating real-time learning capabilities to dynamically adapt to new types of defects as they emerge. Addressing the current limitations by increasing the processing capacity of the model and reducing dependency on lengthy epoch training will be key. Additionally, exploring the integration of the IRP with other industrial monitoring systems could broaden its applicability, ensuring it remains a versatile and effective tool in the rapidly evolving landscape of manufacturing technologies.
6 Conclusions
@@
The experimental validation of the Iterative Refinement Process (IRP) presented in this study underscores its substantial efficacy in enhancing anomaly detection across diverse industrial settings. Employing robust statistical measures and sophisticated anomaly scoring mechanisms, the IRP has consistently demonstrated superior performance over traditional methods such as the vanilla model and One Shot Removal (OSR) model, particularly in environments characterized by high noise levels. Through rigorous testing on the KSDD2 and MVTec-AD datasets, the IRP not only achieved high AUROC scores but also exhibited remarkable resilience against variable noise intensities, effectively maintaining its defect detection capabilities even under challenging conditions. The system’s ability to accurately distinguish between defective and non-defective items, even with subtle defect features, highlights its potential as a critical tool for maintaining high standards in industrial quality control. These results validate the effectiveness of the IRP, showcasing its adaptability to different defect types and its robust performance across a spectrum of industrial products. The study’s outcomes suggest that the IRP can greatly enhance the precision and reliability of defect detection systems, ensuring significant improvements in quality assurance processes within manufacturing environments. {credits}
6.0.1 Acknowledgements
@@
[…]
The authors declare that they have no conflict of interest.

### v2

@@ 4.3 Performance on KSDD2 Across Varying Noise Levels
This section presents the evaluation results of the IRP on the KSDD2 dataset. The analysis explores the performance of the IRP model under different noise levels, highlighting its defect detection capabilities in challenging environments. As shown in Fig. 3, the IRP model is compared against two conventional approaches—the traditional vanilla model (DifferNet) and the One Shot Removal (OSR) model proposed in [Aqeel et al., 2024]. Across all tested noise levels, the IRP model consistently outperforms the competitors, demonstrating remarkable resilience by maintaining superior AUROC values, as systematically documented in Table 1.
The IRP model excels particularly at higher noise levels. For example, at 50% noise, the IRP model achieves a significantly higher AUROC of 0.8967 on the KSDD2 dataset, while the vanilla and OSR models show notable declines. This robustness stems from the IRP’s advanced noise-cancellation algorithms, adaptive thresholding, and self-supervised feature extraction, which enable effective handling of heavily corrupted data.
@@
Moreover, the statistical analysis accompanying our findings reinforces the IRP model’s superiority. The standard deviation values reported in Table 1 indicate lower variability in performance across different experimental runs, underscoring the model’s reliability and predictability. Such attributes are indispensable in industrial applications, where the cost of false negatives can be prohibitive, and the ability to detect anomalies reliably is paramount. The comprehensive data presented in the table not only showcases the IRP model’s adaptability and robust performance across diverse and challenging scenarios but also highlights its viability as a potent tool for maintaining stringent quality control standards in manufacturing processes. These findings suggest that the IRP, through its self-supervised learning approach, can be effectively deployed in industrial settings to enhance defect detection where traditional models might falter due to high variability in defect rates and challenging noise conditions.
@@
Category Entity Noise Level
0% 10% 20% 30% 40% 50%
Classes Bottle 0.9386 0.9495 0.9250 0.9200 0.9295 0.9340
Capsule 0.8592 0.8498 0.8288 0.8102 0.8012 0.7872
Cable 0.9288 0.8368 0.8560 0.6756 0.5716 0.4604
Carpet 0.8386 0.7727 0.7222 0.6354 0.5905 0.6206
Leather 0.9722 0.9388 0.9348 0.9266 0.9157 0.8897
MetalNut 0.8880 0.8688 0.8480 0.8464 0.7792 0.8096
Pill 0.8888 0.8580 0.8558 0.8016 0.8096 0.7816
Screw 0.8938 0.8680 0.8524 0.8234 0.7988 0.7436
Tile 0.9723 0.9674 0.9649 0.9656 0.9598 0.9481
Zipper 0.9378 0.8880 0.8504 0.8232 0.8096 0.7340
Models Vanilla 0.8931 0.8455 0.8235 0.7797 0.7331 0.7171
OSR 0.8960 0.8674 0.8327 0.7916 0.7684 0.7338
IRP on MVTec-AD 0.9118 0.8797 0.8638 0.8228 0.7965 0.7708
IRP on KSDD2 0.9403 0.9238 0.9024 0.9109 0.8923 0.8967
Standard Deviation ±0.0246 ±0.0149 ±0.0086 ±0.0518 ±0.0777 ±0.1099
4.4 Performance on MVTec Across Varying Noise Levels
The robustness and adaptability of IRP have been rigorously evaluated across various product classes within the MVTec-AD dataset, including Bottle, Cable, Capsule, Carpet, Leather, MetalNut, Pill, Screw, Tile, and Zipper. Each class, with its distinct defect types and unique challenges, served as an ideal testbed for the IRP. We systematically varied the noise levels in the dataset from 0% to 50% to meticulously assess the IRP’s capabilities under progressively challenging conditions, using the Area Under the Curve of the Receiver Operating Characteristic (AUROC) to quantify defect detection accuracy across different thresholds.
As illustrated in Fig. 4, the IRP model consistently achieves high AUROC values across all classes, demonstrating its robust ability to manage noise effectively. Notably, in classes with complex and variable defect characteristics such as Screw and Cable, the IRP exhibits exceptional resilience, showcasing its advanced anomaly detection capabilities. Conversely, in classes like Capsule and Pill, where defects are more subtle, the IRP effectively distinguished between normal variations and actual defects, underlining its sophisticated feature extraction capabilities.
[Bottle]Refer to caption [Cable]Refer to caption [Capsule]Refer to caption [Carpet]Refer to caption [Leather]Refer to caption [MetalNut]Refer to caption [Pill]Refer to caption [Screw]Refer to caption [Tile]Refer to caption [Zipper]Refer to caption
@@ 4.4 Performance on MVTec Across Varying Noise Levels
This uniform and robust performance across varied classes and noise levels highlights the IRP model’s adaptability and effectiveness in industrial applications, especially in environments with diverse and unpredictable defect rates. The IRP’s ability to maintain high accuracy at elevated noise levels underscores its potential as a dependable tool for quality control in manufacturing environments. Furthermore, Fig. 3 provides an overarching view of the IRP’s performance across the MVTec-AD dataset. This visual representation emphasizes the IRP’s superior performance relative to the vanilla and OSR models, mainly as it consistently improves defect handling under increasingly challenging conditions. This enhancement affirms the IRP’s role as a robust and reliable quality control instrument in complex manufacturing landscapes, where defect rates fluctuate significantly. Table 1 complements these insights by detailing the AUROC scores at various noise levels for the MVTec-AD and KSDD2 datasets, including measures of error magnitude. This detailed presentation not only confirms the performance stability of the IRP across different settings but also highlights the model’s reliability, evidenced by the low standard deviation in its performance metrics.
4.5 Quantitative Analysis of Removed Samples
This section presents a detailed quantitative analysis of the samples removed during the defect detection process for both the KSDD2 and MVTec-AD datasets. This analysis is segmented by product categories and noise levels, providing insights into the effectiveness of the Iterative Refinement Process (IRP) in improving defect detection accuracy. Table 2 shows the number of good and bad samples removed at various noise levels (0% to 50%) across different product categories within the KSDD2 and MVTec-AD datasets. The categories include Cable, Capsule, Metalnut, Pill, Screw, and Zipper. Each entry in the table specifies the count of good (non-defective) and bad (defective) samples removed during the training and validation phases. The accompanying image Fig. 5 illustrates the samples that were removed during iterative refinement, showcasing both defective samples, which exhibit noticeable defects and noise patterns, and some good samples that were also deleted as part of the refinement process.
The table indicates that at 0% noise, only good samples are removed. This might seem counterintuitive, but it can be explained by the fact that these good samples are likely close to the decision boundary, making them appear anomalous during the initial iterations of the IRP. Removing these borderline good samples helps refine the model by reducing potential noise and improving its generalization capability. This phenomenon demonstrates the IRP’s ability to fine-tune the model even when no bad samples are present. As noise levels increase, the number of bad samples removed rises, which is expected. However, the table also shows that good samples remain removed across all noise levels. This ongoing removal of good samples close to the boundary helps maintain the model’s robustness by continuously refining its decision boundaries.
The quantitative analysis of removed samples underscores the IRP model’s capacity to enhance defect detection accuracy through systematic refinement. By removing samples close to the decision boundary, the IRP improves the overall performance, even in the absence of bad samples. As noise levels increase, the model adeptly identifies and eliminates defective samples while maintaining a balance by removing borderline good samples. This approach fine-tunes the model and ensures its robustness and reliability in real-world industrial applications.
Figure 5: A visual representation of good (outlined in green) and defective (outlined in red) samples under different noise levels. Defects are marked with red circles, illustrating the distribution of flaws within the samples.
@@ 5 DISCUSSION
The robust performance of the Self-Supervised Iterative Refinement Process (IRP) across varying noise levels underscores its potential utility in industrial settings, where defect patterns may be unpredictable. The model’s ability to maintain high AUROC values under increasing noise conditions illustrates its capacity to effectively distinguish subtle defect features from normal variability, which is an essential attribute for critical quality control environments.
@@
While the model demonstrates strong performance, there is scope to enhance its operational efficiency. The current training setup, designed for moderate scale and tailored to specific experimental conditions, serves well under controlled scenarios. However, as we aim to deploy the model in more dynamic industrial settings, optimizing its capacity to handle larger datasets and faster processing cycles will be crucial. These enhancements are intended to reduce the time required to reach model convergence and increase image processing throughput, ensuring the model can meet the demands of real-time defect detection. Future developments could include refining the training algorithms to accelerate convergence without compromising accuracy, and expanding the system’s architecture to support more simultaneous operations. These improvements will be critical for deploying the model in environments where rapid decision-making is essential. Furthermore, the detailed analysis of different classes within the MVTec-AD dataset highlights the IRP’s adaptability to diverse manufacturing scenarios. The pronounced resilience in classes with intricate defect patterns, such as Screws and Cables, underscores the model’s sophisticated feature recognition capabilities, which are likely bolstered by advanced data preprocessing and anomaly scoring mechanisms. Further research could focus on optimizing the model to enhance detection accuracy in scenarios with extremely subtle defect signals, such as those found in high-grade pharmaceutical manufacturing.
@@
Future research efforts might also refine the probabilistic models used within the IRP, enhancing their sensitivity to minor anomalies and incorporating real-time learning capabilities to adapt to new types of defects as they emerge dynamically. Addressing the current limitations by increasing the processing capacity of the model and reducing dependency on lengthy epoch training will be critical. Additionally, exploring the integration of the IRP with other industrial monitoring systems could broaden its applicability, ensuring it remains a versatile and effective tool in the rapidly evolving landscape of manufacturing technologies.
6 CONCLUSION
@@
The experimental validation of the Self-Supervised Iterative Refinement Process (IRP) presented in this study underscores its substantial efficacy in enhancing anomaly detection across diverse industrial settings. Employing robust statistical measures and sophisticated anomaly scoring mechanisms, the IRP has consistently demonstrated superior performance over traditional methods such as the vanilla model and One Shot Removal (OSR) model, particularly in environments characterized by high noise levels. Through rigorous testing on the KSDD2 and MVTec-AD datasets, the IRP not only achieved high AUROC scores but also exhibited remarkable resilience against variable noise intensities, effectively maintaining its defect detection capabilities even under challenging conditions. The system’s ability to accurately distinguish between defective and non-defective items, even with subtle defect features, highlights its potential as a critical tool for maintaining high standards in industrial quality control. These results validate the effectiveness of the IRP, showcasing its adaptability to different defect types and its robust performance across a spectrum of industrial products. The study’s outcomes suggest that the IRP can greatly enhance the precision and reliability of defect detection systems, ensuring significant improvements in quality assurance processes within manufacturing environments.
ACKNOWLEDGEMENTS